Circular No. 7400

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 M. L. McCollough, Universities Space Research Association and Mar-

shall Space Flight Center (MSFC); C. A. Wilson, NASA/MSFC; and X. Sun, University of Alabama in Huntsville, report for the Compton Gamma Ray Observatory (CGRO) BATSE team: “XTE J1550–564 (IAUC 7008, 7010) was detected in outburst with BATSE in the band 20–300 keV on Apr. 6 at a flux of 0.5 Crab (1.0 × 10^{−8} erg cm^{−2} s^{−1}) and had risen to 0.6 Crab (1.2 × 10^{−8} erg cm^{−2} s^{−1}) on Apr. 8. XTE J1550–564 is clearly being detected up to 300 keV. The date that this outburst began is uncertain because XTE J1550–564 was unobservable with CGRO/BATSE from Mar. 26 to Apr. 5. The CGRO/BATSE light curve shows evidence for variability on time scales of less than a day. The energy spectrum for XTE J1550–564 can be fitted by a bremsstrahlung spectrum with a temperature of 101 ± 8 keV or a broken power law with a break energy of 76 ± 25 keV, a lower energy photon index of −1.4 ± 0.4, and a high energy photon index of −2.4 ± 0.4.”

R. Jain and C. Bailyn, Yale University, report YALO consortium observations of the optical counterpart of XTE J1550–564 using the Yale 1-m telescope (+ ANDICAM instrument) at Cerro Tololo Interamerican Observatory: “XTE J1550–564 has entered a new optical outburst. By Mar. 24.4 UT, the object was at V = 21.0, significantly brighter than any of our 12 observations obtained during the period Feb. 28–Mar. 17 (in all of which the source was fainter than V = 21.4). On Mar. 29.3, we find V = 19.07 ± 0.03, and on Apr. 6.3, V = 18.24 ± 0.03. The rate of brightening has decreased in the past two days to V = 18.09 ± 0.02 and 18.06 ± 0.02 on Apr. 9.4 and 10.4, respectively. Further observations at other wavelengths are encouraged.”

SUPERNOVA 1999el IN NGC 6951 CCD magnitudes, unfiltered unless otherwise noted: 1999 Nov. 5.503

UT, 14.5 (S. Yoshida and K. Kadota, Ageo, Saitama, Japan); 13.484, 14.8 (Yoshida and Kadota); 2000 Jan. 28.740, R = 17.3 ± 0.2 (L. Kiss, K. Sárneczky, and K. Sziládi, Piszkéstető, Hungary); 28.816, 16.7 (Yoshida and Kadota); Mar. 16.11, 17.7 ± 0.2 (Sárneczky, Sziládi, and Kiss).

2000 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green